given population 2,7,9 what is the population variance



1
Expert's answer
2022-03-15T03:29:47-0400

Mean: "\mu=\frac{2+7+9}{3}=6."

Variance: "\sigma^2=\frac{(2-6)^2+(7-6)^2+(9-6)^2}{3}=\frac{26}{3}."